# CMB line-of-sight integrators for nearly-isotropic cosmological models

### v1

@@ IV Anilos and Aniclass
Input Bianchi parameters ($\sqrt{h},\Omega^{0}_{K}$) FLRW parameters ($H_{0}$, $\Omega_{m}$, etc.) Precision parameters ($\ell_{\text{max}},\ell_{\text{cutoff}}$, $\tau_{\text{reio}}$) Pert. type (i.e., $|m|$) and vector initial conditions (isocurvature or octopole) Bianchi model $\bm{\leftrightarrow}$ Homog. limit Computes the wavenumber $\nu_{m}$ from Eq. (42) Background Calls CLASS to obtain background parameters as a function of time (${\cal H}$, $a$, $\chi$, $\tau$, etc.) Boltzmann-Bianchi Hierarchy Computes ${}_{s}\kappa^{m}_{\ell}$ and $\zeta^{m}_{\ell}$ Solves (17)-(20) up to $\ell_{\text{max}}$ Radial functions Computes electric and magnetic radial functions Sources Computes the gravitational and collisional sources Line-of-sight integrals Integrates Eqs. (32)-(34). CMB multipolar coefficients $a_{\ell m}^{T}$, $a_{\ell m}^{E}$ and $a_{\ell m}^{B}$
@@
Both codes give the resulting anisotropies for vectors and/or tensor modes having the initial conditions described above. Since scalar modes do not possess regular solutions, they are not implemented. With the exception of model I, which does not possess regular solutions, the codes give the resulting anisotropies from all models in Table 1. In practice, however, we only need to implement models VIIh and IX, since models VII0 and V result from the former in the limits $\ell_{c}\rightarrow\infty$ and $\ell_{s}\rightarrow\infty$, respectively. Model IX is implemented separately, since in this case the coefficient $\zeta^{m}_{\ell}$ achieves a very simple form, implying that only modes with $\ell=|m|=2$ are excited, thus greatly simplifying the hierarchy and radial functions (see Eqs. (6.9) and (6.10) in [29]). Thus, the central core of AniLoS and AniCLASS is devoted to solving model VIIh. In what follows we will detail the structure of AniLoS, while stressing some important differences with AniCLASS.
@@
This module also calls CLASS in order to compute the FLRW background quantities as functions of conformal time, such as ${\cal H}$, $a$, $\chi$ and $\tau$. In AniLoS this is obtained by calling CLASSPython wrapper classy, while in AniCLASS these are obtained directly from the main code. The background quantities consist of arrays in conformal time, with starting and ending values determined by CLASS’s native limits [49].
@@
and, since the source terms do not depend on $\ell$, they could be stored as one-dimensional arrays as $[S^{(m)}_{j}(\eta_{\text{ini}}),\cdots,S^{(m)}_{j}(\eta_{0})]$. In practice, it is more efficient to use NumPy’s function tile to convert it into two-dimensional arrays,
@@
with similar formulas for the polarization multipoles (with an additional $-1$ factor for the $B$ mode). These are the main output of our codes. The coefficients of both the total angular momentum basis and the usual spherical harmonic basis are available in AniCLASS (that is, the user can choose to output $a^{T}_{\ell m}$ or $\Theta^{(m)}_{\ell}$, for example). In AniLoS, only the former is implemented.
Note that, since nearly anisotropic models can only excite coefficients with $|m|=1,2$, many of the resulting coefficients will be zero. Thus, both in AniLoS and AniCLASS, the user can choose between a dense output (i.e., with only non-zero coefficients) or the full output in healpy format (a Jupyter notebook with examples can be found in [35]). However, in AniCLASS, the full output is not available in the Python wrapper. Another important aspect of both codes is that the equations implemented are specialized to the (fixed) polariation basis $(\bm{e}^{(+)},\bm{e}^{(-)},\bm{e}_{3})$. When doing forecasts or real-data search for anisotropies, one should also allow for three Euler angles corresponding to the orientation of the (unknown) axes of anisotropies. In practice, this can be done by rotating the output $a_{\ell m}$s with HEALPix/healpy rotation routines.
@@ V Examples
Figure 3: Evolution of $\beta^{\prime}_{(m)}$ and the neutrino quadrupole ${\cal N}^{(m)}_{2}$ for vector (left) and tensor (right) modes, and both for real (continuous) and imaginary (dashed) parts. For this plot we have used: $\Omega_{K}^{0}=0.1$ (VIIh and V), $10^{-5}$ (VII0), and $-0.1$ (IX); $\sqrt{h}=0.01$ (VIIh), $10^{4}$ (V), and $10^{-3}$ (VII0). For the vector modes, the neutrino isocurvature quadrupole initial condition was used. Recall that model IX has no vector modes, and is thus shown only on the right panel.
Let us now illustrate our codes with a few examples. Figure 3 illustrates the evolution of vector and tensor modes of the shear and neutrino quadrupole. Note how the tensor component of the shear is initially constant despite the small values of anisotropic stress (right panel). Conversely, the vector component of the shear is only initially constant because of the intensity of the anisotropic stress. Figures 4 to 7 showcase the expected temperature and polarization CMB anisotropies in all nearly-isotropic models (except model I), both for regular tensor modes and the two possible regular vector modes we considered. In all these figures, the $z$-axis was chosen to point along the normal to the screen, so as to help with the visualization of the anisotropic features.
Models VII0 and VIIh are the only cases where spatial translations are accompanied by corkscrew rotations. Consequently, these models display a clear spiraling pattern seen in Figures 4 and 5 (see also [7]). Notice how the spiralling patterns for the vector modes are stronger than those found in the tensor modes (2nd and 3rd rows in both figures). This is because the typical spiral scale is given by $2\pi\ell_{s}/|m|$ in these models. However, from the observational point of view, the most important feature of these models lies in the multipolar structure, which goes beyond a simple quadrupole $\ell=2$. This follows from the non-trivial structure of the $\zeta^{m}_{\ell}$ coefficients entering the hierarchy (see Table 1), which in turn reflects the non-trivial kinematics of light in these models. This is in sharp contrast to the anisotropies of models V and IX shown in Figures 6 and 7, which can only display quadrupolar anisotropies. Models V and VIIh also display a focusing of the quadrupole pattern along the $z$-axis, which is a typical signature of hyperbolic spaces [50].
@@ VI Conclusions
Upcoming cosmological observations hold the potential to deepen our understanding of the universe, including a better assessment of its geometry at horizon scales. This task will not only benefit from new tools allowing an efficient comparison between data and models, but also an integration of existing tools. In this work, we have introduced two numerical softwares, AniLoS and AniCLASS, allowing for a unified treatment of stochastic and deterministic CMB anisotropies. In particular, when the deterministic component arises from nearly-isotropic cosmological models, it was shown in [29] that the resulting anisotropies can be obtained from standard line-of-sight CMB codes with minimal modifications.
Although both softwares perform the same tasks, they have different use cases. AniLoS is a user-friendly Python package allowing for easy inclusion of new initial conditions and exotic physics, whereas AniCLASS, an extension of CLASS, targets at efficiency and intensive Monte Carlo simulations. Both codes are free to download and use at https://github.com/jgvicente.

### v2

@@ IV Anilos and Aniclass
[…]
@@
Both codes give the resulting anisotropies for vectors and/or tensor modes having the initial conditions described above. Since scalar modes do not possess nontrivial regular solutions, they are not implemented. With the exception of model I (which does not possess regular nontrivial solutions), the codes give the resulting anisotropies from all models in Table 1.
In practice, however, we only need to implement models VIIh and IX, since models VII0 and V result from the former in the limits $\ell_{c}\rightarrow\infty$ and $\ell_{s}\rightarrow\infty$, respectively. Model IX is implemented separately, since in this case the coefficient $\zeta^{m}_{\ell}$ achieves a very simple form, implying that only modes with $\ell=|m|=2$ are excited, thus greatly simplifying the hierarchy and radial functions (see Eqs. (6.9) and (6.10) in [55]). Thus, the central core of AniLoS and AniCLASS is devoted to solving model VIIh. In what follows we will detail the structure of AniLoS, while stressing some important differences with AniCLASS.
@@
[…]
@@
and, since the source terms do not depend on $\ell$, they could be stored as one-dimensional arrays as $S^{(m)}_{j}(\eta_{\text{ini}}),\cdots,S^{(m)}_{j}(\eta_{0})]$. In practice, it is more efficient to use NumPy’s function tile to convert it into two-dimensional arrays,
@@
with similar formulas for the polarization multipoles (with an additional 1-1 factor for the $B$ mode). These are the main output of our codes. The coefficients of both the total angular momentum basis and the usual spherical harmonic basis are available in AniCLASS (that is, the user can choose to output $a^{T}_{\ell m}$ or $\Theta^{(m)}_{\ell}$, for example). In AniLoS, only the former is implemented.
Note that, since nearly anisotropic models can only excite coefficients with $|m|=1,2$, many of the resulting coefficients will be zero. Thus, both in AniLoS and AniCLASS, the user can choose between a dense output (i.e., with only non-zero coefficients) or the full output in healpy format (a Jupyter notebook with examples can be found in [72]). However, in AniCLASS, the full output is not available in the Python wrapper. Another important aspect of both codes is that the equations implemented are specialized to the (fixed) polariation basis $(\boldsymbol{e}^{(+)},\boldsymbol{e}^{(-)},\boldsymbol{e}_{3})$. When doing forecasts or real-data search for anisotropies, one should also allow for three Euler angles corresponding to the orientation of the (unknown) axes of anisotropies. In practice, this can be done by rotating the output $a_{\ell m}$s with HEALPix/healpy rotation routines.
@@ V Examples
Figure 3: Left: evolution of $\beta^{\prime}_{(1)}$ and the neutrino quadrupole ${\cal N}^{(1)}_{2}$ for vector modes. Right: evolution of $\beta_{(2)}$ and the neutrino quadrupole ${\cal N}^{(2)}_{2}$ for tensor modes. Both panels show the real (continuous) and imaginary (dashed) parts of the solution. For this plot we have used: $\Omega_{K}^{0}=0.1$ (VIIh and V), $10^{-5}$ (VII0), and 0.1-0.1 (IX); $\sqrt{h}=0.01$ (VIIh), $10^{4}$ (V), and $10^{-3}$ (VII0). For the vector modes, the neutrino isocurvature quadrupole initial condition was used. Recall that model IX has no vector modes, and is thus shown only on the right panel.
Let us now illustrate our codes with a few examples. Figure 3 illustrates the evolution of vector and tensor modes of the shear and neutrino quadrupole. Note how the tensor component of the shear is initially constant despite the small values of anisotropic stress (right panel). Conversely, the vector component of the shear is only initially constant because of the intensity of the anisotropic stress. The right panel of Figure 3 also displays the trivial tensor solution for model V, normalized as $\beta_{(\pm 2)}\equiv 1$, and which follows as a limit of the tensor solution of model VIIh, as explained above. Figures 4 to 7 showcase the expected temperature and polarization CMB anisotropies in all nearly-isotropic models (except model I), both for regular tensor modes and the two possible regular vector modes we considered. In all these figures, the $z$-axis was chosen to point along the normal to the screen, so as to help with the visualization of the anisotropic features.
Models VII0 and VIIh are the only cases where spatial translations are accompanied by corkscrew rotations. Consequently, these models display a clear spiraling pattern seen in Figures 4 and 5 (see also [61]). Notice how the spiralling patterns for the vector modes are stronger than those found in the tensor modes (2nd and 3rd rows in both figures). This is because the typical spiral scale is given by $2\pi\ell_{s}/|m|$ in these models. [^6] The best available constraints on this scale come from the Wilkinson Microwave Anisotropy Probe (WMAP) data, and give $x\equiv\ell_{s}/H_{0}^{-1}\approx 1.2$ [65]. However, the posterior distribution of $x$ is very broad (see Figure 5 of Ref. [65]), so that current constraints on $\ell_{s}$ are weak. However, from the observational point of view, the most important feature of these models lies in the multipolar structure, which goes beyond a simple $\ell=2$. This follows from the non-trivial structure of the $\zeta^{m}_{\ell}$ coefficients entering the hierarchy (see Table 1), which in turn reflects the non-trivial kinematics of light in these models. This is in sharp contrast to the anisotropies of models V and IX shown in Figures 6 and 7, which can only display quadrupolar anisotropies. Models V and VIIh also display a focusing of the quadrupole pattern along the $z$-axis, which is a typical signature of hyperbolic spaces [14].
@@ VI Conclusions
Upcoming cosmological observations, particularly those probing near-horizon scales — such as LiteBIRD [9] — have the potential to significantly improve our understanding of the Universe, including more robust constraints on its geometry at horizon scales. This task will not only benefit from new tools allowing an efficient comparison between data and models, but also an integration of existing tools. In this work, we have introduced two numerical softwares, AniLoS and AniCLASS, allowing for a unified treatment of stochastic and deterministic CMB anisotropies. In particular, when the deterministic component arises from nearly-isotropic cosmological models, it was shown in [55] that the resulting anisotropies can be obtained from standard line-of-sight CMB codes with minimal modifications.
Although both softwares perform the same tasks, they have different use cases. AniLoS is a user-friendly Python package allowing for easy inclusion of new initial conditions and exotic physics, whereas AniCLASS, an extension of CLASS, targets at efficiency and intensive Monte Carlo simulations. Both codes are free to download and use at [72].
@@
The possibility of including non-trivial initial conditions may be relevant for investigating the origin of low-$\ell$ CMB anomalies. As illustrated in the figures of Section V, a quadrupolar modulation of temperature and polarization anisotropies is an ubiquitous signature of the shear. At the same time, the low value of the CMB quadrupole [19, 20, 5] and its alignment with the octopole [25, 67] are one of the most robust statistical anomalies, the latter persisting even after correction for the frequency-dependent kinetic Doppler quadrupole [51]. Although the amplitude of the shear is tightly constrained by Planck [66], many of these anomalies could be associated with off-diagonal correlations between $a_{\ell m}$’s arising from a possible breakdown of statistical isotropy, which could in turn evade such constraints. Whether these anomalies are physical and, if so, of a deterministic or stochastic origin, is yet not known.